### v1

# Probing the cosmic sterile-neutrino background with IceCube

Abstract
In this paper, we take a close look at the interaction between the TeV–PeV energy astrophysical neutrinos and a hypothetical cosmic sterile-neutrino background. These interactions yield absorption features, also called “dips”, in the astrophysical neutrino spectrum, which are studied using the deposited energy distribution of high-energy starting events (HESE) in the IceCube detector. We improve upon the previous analysis by including the effects of regeneration and a realistic source distribution on the propagation of astrophysical neutrinos. We use the latest 7.5-year HESE dataset and include the observation of Glashow resonance in our analysis. We evaluate the impact of these dips on the inferred spectral index and overall normalization of the astrophysical neutrinos. We find a mild preference for dips in the 300–800 TeV range, and the best-fit parameters for the mass of sterile-neutrino and the mediator are 0.5 eV and 23 MeV, respectively. We find that the inclusion of these absorption features lowers the spectral index of astrophysical neutrinos to $2.60^{+0.19}_{-0.16}$. The lower spectral index can reduce the disagreement with the Northern Tracks sample but requires dedicated analysis. We also forecast the event spectrum for IceCube-Gen2 for the two different fits.
I Introduction
The IceCube neutrino observatory has detected neutrinos of astrophysical origin in the TeV–PeV range in several detection channels. The latest 7.5 years (2635 days) high energy starting events (HESE) sample consists of 102 events, of which 60 events have reconstructed deposited energy above 60 TeV Abbasi et al. (2021). The IceCube experiment has also observed a Glashow Resonance (GR) event Glashow (1960) in the 4.6 years PeV energy partially contained events (PEPE) sample Aartsen et al. (2021a). IceCube also maintains a catalog of through-going muon-neutrinos from the northern hemisphere (henceforth called Northern Tracks). A large fraction of the Northern Tracks are below 100 TeV and originate via atmospheric neutrinos. The highest energy events from this sample have astrophysical origin.
It is abundantly clear that these neutrinos have an extra-galactic origin and are likely to originate in sources capable of accelerating cosmic rays to EeV-scale energies. The neutrinos are produced in subsequent interactions of these ultra-high energy cosmic rays with ambient gas or radiation. At the time of writing, only a handful of such sources have been confirmed Aartsen et al. (2018a, b, 2020); Abbasi et al. (2022a). If the neutrinos are produced via interactions with ambient gas, i.e. in proton-proton ($pp$) interactions, the neutrinos are expected to have a power-law spectrum Margolis et al. (1978); Stecker (1979); Kelner et al. (2006). In the baseline scenario, the neutrinos oscillate and reach flavor equilibrium before they arrive to the Earth Lipari et al. (2007); Choubey and Rodejohann (2009); Mena et al. (2014, 2014); Aartsen et al. (2015). One usually assumes that an isotropic flux of astrophysical neutrinos is incident on the Earth with a total differential all-flavor neutrino-plus-anti-neutrino spectrum given by
$\dfrac{d\Phi_{6\nu}}{dE_{\nu}}=\Phi_{\texttt{astro}}\,\left(\frac{E_{\nu}}{E_{0}}\right)^{-\gamma_{\texttt{astro}}}\,\cdot 10^{-18}\,\text{GeV}^{-1}\,\text{cm}^{-2}\,\text{s}^{-1}\,\text{sr}^{-1},$ (1)
which is characterized by two free parameters: a spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$). The reference energy scale $E_{0}$ is chosen to be 100 TeV. If the neutrinos are produced via photo-hadronic interactions, i.e. in proton-gamma ($p\gamma$) sources, the energy spectrum is expected to be “bump-like” Kelner et al. (2006); Mucke et al. (2000); Hummer et al. (2010); Fiorillo and Bustamante (2023). Realistically, both $pp$ and $p\gamma$ sources are expected to contribute. In this work, we will assume that the $pp$ sources dominate and the diffuse flux of astrophysical neutrinos is described by a single power-law (SPL), as in Equation 1.
Are the events observed by IceCube consistent with the assumption of SPL? To a large extent, yes. However, there are some tantalizing discrepancies that have motivated a large body of work in the last decade. Some of the explanations are based on dark matter Feldstein et al. (2013); Esmaili and Serpico (2013); Bai et al. (2016); Fong et al. (2015); Rott et al. (2015); Esmaili et al. (2014); Murase et al. (2015); Anchordoqui et al. (2015); Bhattacharya et al. (2017a, b); Hiroshima et al. (2018); Chakravarty et al. (2020), resonances beyond standard model Anchordoqui et al. (2006); Barger and Keung (2013); Dutta et al. (2015); Dey and Mohanty (2016); Dev et al. (2016); Mileo et al. (2016); Chauhan et al. (2018); Dey et al. (2018); Bečirević et al. (2018); Babu et al. (2020); Dey et al. (2021), or non-standard/secret neutrino interactions Ng and Beacom (2014); Ioka and Murase (2014); Ibe and Kaneta (2014); Blum et al. (2014); Araki et al. (2015); Cherry et al. (2014); Kamada and Yu (2015); DiFranzo and Hooper (2015); Araki et al. (2016); Rasmussen et al. (2017); Jeong et al. (2018); Mohanty et al. (2019); Pandey et al. (2019a, b); Bustamante et al. (2020); Creque-Sarbinowski et al. (2021); Mazumdar et al. (2022); Fiorillo et al. (2020a, b); Carpio et al. (2023); Esteban et al. (2021). One such discrepancy is the inferred value of the spectral index in two independent datasets,
[…]
We begin with the ansatz that including “dips” in the neutrino flux will result in a lower spectral index from the HESE sample and thus relax the tension. Such dips are a consequence of interaction between astrophysical neutrinos and cosmic background. If the interaction is through a s-channel process, then the absorption feature is narrowly located around the resonance energy $E_{\nu}^{\rm res}=M_{X}^{2}/2m_{b}$, where $M_{X}$ is the mass of the mediator and $m_{b}$ is the mass of the background particle. If the interaction is t-channel, then a broad spectrum feature is expected Mohanty et al. (2019). In Ref. Esteban et al. (2021), the effect of non-resonant contribution to the process was found to be significant for larger couplings, and we have included them in our analysis. With only standard model interactions, the dip will occur for neutrinos with energy EeV Weiler (1982); Brdar et al. (2022). However, if there are non-standard/secret interactions mediated by a MeV-scale boson, then the dips can occur in the TeV–PeV range Ng and Beacom (2014). In this paper, we study the scenario proposed in Ref. Chauhan and Mohanty (2018) where the dip is caused by interactions between astrophysical neutrinos and a cosmic sterile-neutrino background.
Figure 1: The test statistic (see Equation 5) for the single power-law fit of 7.5 years HESE sample Abbasi et al. (2021) is shown in blue. In orange, we show the test statistic after including absorption features and Glashow resonance. The horizontal grey dashed lines are used to denote the 68% C.L. and 90% C.L. intervals. We have shown the results from the 11-year Northern Tracks sample Stettner (2020) in green.
The light sterile neutrino is a well-motivated extension of the standard model and has been the subject of investigation in a number of experiments. Currently, there is evidence in support of  Anselmann et al. (1995); Hampel et al. (1998); Aguilar-Arevalo et al. (2009, 2010, 2001); Mention et al. (2011); Aguilar-Arevalo et al. (2018); Serebrov et al. (2021) as well as against Abratenko et al. (2023) the sterile neutrino. For a recent review, we refer to Ref. Dasgupta and Kopp (2021). The presence of a fully thermalized sterile neutrino is in conflict with the measurement of $N_{\text{eff}}$ from cosmological probes Hamann et al. (2011); Cooke et al. (2018); Aghanim et al. (2020). If the sterile neutrino has self-interactions via a MeV scale vector mediator, their recoupling can be delayed, and the $N_{\text{eff}}$ constraints can be avoided Hannestad et al. (2014); Dasgupta and Kopp (2014). The case for pseudo-scalar mediator is studied in Ref. Archidiacono et al. (2015). Despite the early success, the self-interacting sterile neutrino scenario has been tested against many cosmological observations Saviano et al. (2014); Mirizzi et al. (2015); Forastieri et al. (2017); Choi et al. (2018); Song et al. (2018). It was shown in Ref. Chu et al. (2018) that the basic scenario is in conflict with bounds on the sum of neutrino mass, and some resolutions have been provided as well. One must note that these results depend on the choice of background cosmology as $\Lambda$CDM, and the bounds will change if other scenarios are considered Roy Choudhury and Naskar (2019); Roy Choudhury and Choubey (2018); Roy Choudhury and Hannestad (2020). Recent analyses of large-scale structure data have also indicated the preference for self-interacting neutrinos He et al. (2024); Camarena et al. (2023). With this as motivation, we ask if the dip due to a cosmic sterile-neutrino background can be probed in IceCube and whether the inclusion of these features leads to an appreciable decrease in the inferred spectral index. The main result of this paper is shown in Figure 1. We find that the systematic inclusion of the dips results in,
[…]
which is in better agreement with the Northern Tracks sample. As discussed below, the errors are underestimated due to the limitations of our analysis. For a quantitative assessment of the relaxation in the tension between the two datasets, one must evaluate the impact of dips in the spectrum on the Northern Tracks sample. However, this exercise is beyond the scope of our paper. We have shown that an ‘SPL with dips’ is a viable fit to the present HESE sample and results in a lower spectral index than an SPL-only explanation. One must note that our conclusion also extends to other origins of the dip(s).
The paper is organized as follows. In Section II, we discuss the model, propagation of astrophysical neutrinos, our estimate of the HESE event rate, and the likelihood. In Section III, we provide the main results and end with a short summary in Section IV.
II Model description and methods
In this work, we follow the model outlined in Ref. Dasgupta and Kopp (2014), which is briefly described below. The sterile neutrinos are assumed to have self-interactions mediated by a new MeV-scale vector boson. The interaction Lagrangian is given by,
$-\mathcal{L}_{\rm int}=g_{X}\bar{\nu}_{s}\gamma_{\mu}\frac{1}{2}(1-\gamma_{5})\nu_{s}X^{\mu}\,,$ (2)
where $X^{\mu}$ is the mediator and $g_{X}$ is the self-interaction coupling constant. In this work, we fix the coupling strength as $g_{X}=0.1$, which is sufficiently large for the re-coupling of sterile neutrinos Chu et al. (2015). The mass-eigenstates and the flavor-eigenstates are related by a $4\times 4$ rotation matrix characterized by six mixing angles. The three mixing angles from the active sector are fixed to their measured values from Ref. Esteban et al. (2020), and the remaining three angles are free parameters of the model. Due to this mixing, all four mass-eigenstates interact via the new mediator, and interactions between propagating mass-eigenstates and background neutrinos are possible Chauhan and Mohanty (2018). In this simplified model, we have five new physics parameters – three mixing angles ($\theta_{14}\,,\theta_{24}\,,\theta_{34}$), mass of the sterile neutrino ($m_{s}$), and mass of the mediator ($M_{X}$). In a UV-complete model, additional degrees of freedom may be required for generating the mixing and mass of the gauge boson and for anomaly cancellation. The alternative model with a pseudoscalar mediator is cosmologically more viable Dasgupta and Kopp (2021), but has additional constraints from meson decays Fiorillo et al. (2020a). We anticipate that similar bounds will exist on vector mediators, but their evaluation is beyond our scope. Our main conclusion, i.e., the spectral index is lowered because of the dips, does not depend on the details of the model causing the dips in the spectrum.
The introduction of this new interaction results in scattering between astrophysical neutrinos and the cosmic background of active and sterile neutrinos. Consequently, we expect dips and bumps in the flux of astrophysical neutrinos, which is otherwise assumed to have a single-power law behavior. The Boltzmann equation for the time evolution of the comoving number density of the astrophysical neutrinos per unit energy ($\tilde{n}_{i}=\partial\Phi_{i}/\partial E_{\nu}$) is given by Ng and Beacom (2014)
$\frac{\partial\tilde{n}_{i}(t,E_{\nu})}{\partial t}$ $=\frac{\partial}{\partial{E_{\nu}}}\left[H(t)\,E_{\nu}\,\tilde{n}_{i}(t,E_{\nu})\right]+\mathcal{L}_{i}(t,E_{\nu})-\tilde{n}_{i}(t,E_{\nu})\sum_{j}n^{t}_{j}\sigma_{ij}(E_{\nu})$
[…]
In this equation, the first term on the right-hand side accounts for the cosmological expansion. The second term, called the source term, accounts for the production of high-energy neutrinos due to astrophysical sources. In this work, we assume that the TeV–PeV neutrinos originate in $pp$-sources and following Ref. Esteban et al. (2021), the source term is taken to be a single power law with normalization proportional to the star-formation rate. The third term accounts for the interaction with background neutrinos, and the last term accounts for the regeneration of lower-energy neutrinos after interactions. These equations have been solved using the publicly available code nuSIprop Esteban et al. (2021), which solves the Boltzmann equations for standard three-flavor neutrinos in the presence of non-standard interaction. We have extended the code to include the sterile neutrino. These are improvements on the work done in Ref. Chauhan and Mohanty (2018) where the source term was ad-hoc, and regeneration was ignored.
To estimate the rate of high-energy starting event (HESE) in IceCube, we use the Monte Carlo sample of simulated events [^1] https://github.com/icecube/HESE-7-year-data-release/ from Ref. Abbasi et al. (2021). The input parameters used to calculate the HESE event spectrum are tabulated in Table 1, and a description of these parameters can be found in Ref. Abbasi et al. (2021). Of these, we only consider the astrophysical neutrino spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) as variables and fix other parameters to their typical values. We assume that the new physics does not affect the fluxes of conventional and prompt atmospheric neutrinos, as well as atmospheric muons. A detailed calculation where all parameters are optimized is computationally very expensive and beyond the scope of this work. The event spectrum and the subsequent log-likelihood are obtained by reweighting the Monte Carlo events from astrophysical neutrinos. As we also scan over the mixing angles, the reweighting accounts for the flavor dependence of the interactions.
Table 1: The set of physical and nuisance parameters that determine the HESE event rate are tabulated. We have shown the best-fit parameters for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the values/range of parameters used in this work. For a description of these variables, see TABLE VI.1 in Ref. Abbasi et al. (2021)
Parameter SPL best-fit This Work
$\gamma_{\texttt{astro}}$ 2.874 (2,4)
$\Phi_{\texttt{astro}}$ 6.365 (1,10)
$\Phi_{\texttt{conv}}$ 1.006 1.000
$\Phi_{\texttt{prompt}}$ 0.0 0.0
$\Phi_{\texttt{muon}}$ 1.187 1.00
$\Delta\gamma$0.053 -0.050
$\nu/\overline{\nu}$ 0.998 1.000
$a_{s}$ 1.001 1.000
$\varepsilon_{\rm dom}$ 0.952 0.990
$\varepsilon_{\rm h.o.}$0.055 0.000
$K/\pi$ 1.000 1.000
We follow the analysis of Ref. Abbasi et al. (2021) closely and present our results in the frequentist statistical methodology. We report the new physics parameters that are most likely to explain the HESE data and compare it with the single power-law. The best-fit parameters and the errors are obtained using the profile likelihood,
~$\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})=\max_{\bm{\eta}}\mathcal{L}(\bm{\theta,\eta})\cdot\Pi(\bm{\theta,\eta})$ (4)
where $\bm{\theta}$ is the set of parameters of interest and $\bm{\eta}$ is the set of nuisance parameters. The likelihood function used is the effective likelihood defined in Ref. Argüelles et al. (2019). The constraint $\Pi(\bm{\theta,\eta})$ is constant for parameters for which no external data is available. We have used the Python implementation of SLSQP algorithm to minimize the negative of the log of the function and thus obtain the best-fit point $\hat{\bm{\theta}}$. The model parameter test statistic is defined as,
$\text{TS}(\bm{\theta})=-2\,\log\,\left(\dfrac{\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})}{\tilde{\mathcal{L}}^{\texttt{profile}}(\hat{\bm{\theta}})}\right).$ (5)
The results are presented using Wilks’ theorem with appropriate degrees of freedom Wilks (1938). To validate our method, we first evaluate the 68 % C.L. and 90 % C.L. intervals for single-power law and compare them with the contours in Ref. Abbasi et al. (2021). We only scan over the astrophysical flux parameters and keep other parameters fixed to their best-fit values. We find that on expected lines, the best-fit point agrees very well, and the errors are underestimated.
The HESE sample only includes events for which the interaction vertex is inside the restricted fiducial volume. To enhance the sensitivity to Glashow resonance (GR) events, a larger fiducial volume is required, which is achieved by looking at PeV energy partially contained events (PEPE). As the Monte Carlo event sample contains the GR, we use a simple rescaling for exposure to estimate the GR event rate. We find that our estimates agree with the Extended Data Table 1 in Ref. Aartsen et al. (2021a). The likelihood function, i.e., $\mathcal{L}(\bm{\theta,\eta})$ in Equation 4, is modified to include GR event for which we assume a simple Poisson likelihood.
III Searching for dips in spectrum
In this section, we look at the range of new physics parameters that lead to significant dips in the HESE sample and evaluate their impact on the spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) of the diffuse astrophysical flux. In our analysis, we fix the values of the parameters related to the detector efficiency, cosmic ray flux, and atmospheric neutrinos (see Table 1). As a result of this approximation, the errors in this paper are underestimated. A detailed scan of all parameters is computationally intensive and beyond the scope of this work. The parameters of interest are,
$\bm{\theta}=\{\gamma_{\texttt{astro}},\,\Phi_{\texttt{astro}},\,M_{X},\,m_{s},\,\theta_{14},\,\theta_{24},\,\theta_{34}\}.$ (6)
The best-fit values of these parameters from our analysis is tabulated in Table 2. We fix the mixing angles to their best-fit values for the rest of the analysis. This reduces the computation time significantly and allows us to scan over relevant parameter space only. This approximation underestimates the 90% C.L. region, and the 68 % C.L. region is mostly unaffected.
Table 2: The best-fit values of the parameters of interest for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the result from our analysis is tabulated.
Parameter SPL best-fit This Work
$\gamma_{\texttt{astro}}$ 2.874 2.613
$\Phi_{\texttt{astro}}$ 6.365 5.461
$M_{X}$ 22.96 MeV
$m_{s}$ 0.50 eV
$\theta_{14}$ 0.52
$\theta_{24}$ 0.53
$\theta_{34}$ 0.41
Figure 2: Scan of the profile likelihood (see Equation 5) for the two new physics parameters: mass of the sterile neutrino ($m_{s}$) and mass of the vector boson mediator ($M_{X}$), is shown in orange. The darker (lighter) region denotes the 68% C.L. (90 % C.L.) interval. Note that for each point, the mixing angles are fixed to their best-fit values, and all other parameters are optimized. We also show two approximate curves that result in dips at 300 TeV and 800 TeV due to interaction with a cosmic sterile neutrino background. The region labeled as NSI (pink) indicated the parameter space for which dips occur due to interaction with the active neutrino background only and the dip due to sterile neutrino is outside the HESE analysis bound.
We first look at the space of the new physics parameters – sterile neutrino mass ($m_{s}$) and the vector mediator mass ($M_{X}$), and present our results in Figure 2. As expected [^2] The Figure 2 should be compared with Fig. 2 in Chauhan and Mohanty (2018), we find a large degenerate parameter space that results in dips in the event spectrum. We also find an island for $M_{X}=5-10$ MeV, which is largely independent of the sterile neutrino mass. In this region, the dip due to sterile neutrino is below the 60 TeV threshold of the HESE sample, and there is a dip between 300-800 TeV due to the heavier active neutrino. This is similar to the case of neutrino secret interaction ($\nu$SI), which is considered in Ref. Ng and Beacom (2014). We also highlight the $m_{s}=2.6$ eV that is hinted in recent results Serebrov et al. (2021); Abbasi et al. (2022b).
Figure 3: Scan of the profile likelihood for the two new physics parameters: astrophysical normalization and spectral index. The darker (lighter) regions show the 68 % C.L. (90 % C.L.) intervals. The region in blue is obtained using single power law, and the region in orange is obtained by including Glashow Resonance and dips due to sterile neutrino background. Note that other physical parameters are optimized for each scan point. As we do not optimize over all of the parameters (see text for details), the intervals are slightly smaller than the ones reported in Ref. Abbasi et al. (2021).
Figure 4: Left: We show the best-fit SPL flux from Ref. Abbasi et al. (2021) (blue) and the best-fit flux from our analysis (orange). The binned reconstructed flux from the HESE sample Abbasi et al. (2021) and PEPE sample is shown in black and red, respectively. Right: The deposited energy spectrum of HESE events (black) is shown with our event prediction using the best-fit parameters for SPL with dips (orange) and SPL-only (blue). The background from atmospheric neutrinos and muons is the same for both curves. The shaded grey region represents the 60 TeV threshold below which the backgrounds dominate the event rate.
We now look at the space of astrophysical neutrino flux parameters – spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$). The scan of the profile likelihood is presented in Figure 3. We find that the inclusion of the dips leads to lower values of the spectral index. The inclusion of the Glashow resonance event also disfavors large values of the spectral index. This is one of the key results of the paper. At present, with only 7.5 years of data, there is a large overlap between the preferred regions for SPL and SPL with dips. However, the trend is clearly noticeable. In the future, with additional data from other neutrino telescopes, we anticipate that the overlap will be resolved. This was recently demonstrated in Ref. Fiorillo and Bustamante (2023), where the authors search for a bump-like feature in the spectrum. We also look at single parameter profile likelihood, and the results have been presented in Figure 1. It is clear that the inclusion of absorption features leads to a lowering of the spectral index and thus reduces the tension in the HESE and Northern Tracks samples. A future systematic study would also include the impact of absorption on the Northern Tracks sample.
In order to highlight the difference between the SPL best-fit and the best-fit parameters for SPL with dips, we show the neutrino flux prediction overlaid with the reconstructed flux from the IceCube collaboration in Figure 4. This plot only indicates of the position of dips in the neutrino spectrum. In our analysis, the likelihood is computed using the data from the deposited energy event spectrum. We have shown the predicted event spectrum from the two sets of best-fit parameters in Figure 4. As mentioned earlier, we assume that the atmospheric neutrinos are unaffected by the new physics, and the background predictions for both fits are the same. In a more detailed analysis, one may consider varying the atmospheric neutrino parameters. As there is a large overlap in the 68% C.L. interval of the two fits, we choose not to show the uncertainty bands in these figures.
Figure 5: Event rate prediction for 25 years of exposure for IceCube-Gen2. The orange curve represents events assuming the best-fit parameters of this work (see Table 2). The blue curve represents events for the SPL best-fit parameters. The blue-shaded region shows 68 % C.L. errors on the SPL event rate assuming Poisson distribution. The grey-shaded region represents the 200 TeV threshold for IceCube-Gen2 Aartsen et al. (2021b).
III.1 Forecast for IceCube-Gen2
IceCube-Gen2 is the planned upgrade of the IceCube neutrino observatory, which is expected to have $\sim$10 times larger exposure to TeV–PeV neutrinos Aartsen et al. (2021b). To estimate the event rate in IceCube-Gen2, we scale the weights according to the increased effective area. The effective area for cascades in IceCube-Gen2 is given in Ref. Aartsen et al. (2021b). However, the effective area for tracks in IceCube-Gen2 is not publicly available at the time of writing this paper. In this work, we will assume that the effective area is the same for cascades and tracks. We have evaluated the event rates using 25 years of IceCube-Gen2 exposure for the best-fit parameters from the SPL analysis and for the best-fit point from this work. The results are presented in Figure 5. The local significance of the dips can be tested with such large exposures. It must be noted that there are other proposed neutrino telescopes that will also be sensitive to the absorption features. A combined analysis will improve the discovery prospects Fiorillo and Bustamante (2023).
IV Summary and outlook
The IceCube neutrino observatory has played a crucial role in our understanding of astrophysical neutrinos in the last decade. It also provides an opportunity to study rare interactions over large distances. In this paper, we have taken a close look at the interaction of the astrophysical neutrinos with a hypothetical sterile neutrino background. The light sterile neutrino is well-motivated from short baseline experiments. However, there is a conflict with standard cosmology, and one of the resolutions is to allow self-interactions via a MeV-scale mediator. As a consequence, one expects the cosmic neutrino background to have a sterile neutrino component as well. The astrophysical neutrinos can interact with the sterile neutrino background, which leads to dips in the TeV–PeV neutrino energy spectrum. If the dips are significant, there is an observable impact on the HESE sample in IceCube.
We have improved on the previous analysis on multiple fronts. We have taken a realistic source distribution and included the effect of regeneration during the propagation. We also account for the recent observation of Glashow resonance in our analysis. To estimate the likelihood, we first compute the event rate and then compare with the observations. We find that the inclusion of these absorption features allows for a lower spectral index ($\gamma_{\texttt{astro}}=2.60^{+0.19}_{-0.16}$) when compared with the single-power law flux ($\gamma_{\texttt{astro}}=2.87^{+0.20}_{-0.19}$). Note that, due to the approximations in our analysis, our errors are underestimated. The lower spectral index is in better agreement with the Northern Tracks sample ($\gamma_{\texttt{astro}}=2.28^{+0.08}_{-0.09}$). However, a detailed study that accounts for the effect of these absorption features on the Northern Tracks sample is essential and left as future work. Currently, with only 7.5 years of HESE data, we find that there is a large overlap of parameter space, and one cannot statistically infer the existence of such dips. In the future, with a much larger neutrino telescope in the form of IceCube-Gen2 and perhaps other IceCube-scale experiments, the absorption features will be strongly tested.
Acknowledgments
The authors would like to thank Austin Schneider for helping with the HESE Monte Carlo simulation and Ivan Esteban for helping with nuSIprop. The authors would also like to thank Subhendra Mohanty and Ivan Esteban for their comments and suggestions on the manuscript. PP thanks Ranjan Laha for useful discussion. BC is supported by an OPERA grant and seed grant NFSG/PIL/2023/P3821 from BITS Pilani. PP acknowledges the support from the National Post-Doctoral Fellowship by the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (SERB-PDF/2023/002356) and the IOE-IISc fellowship program.

### v2

@@
 We find that the inclusion of these absorption features lowers the spectral index of astrophysical neutrinos to $2.60^{+0.19}_{-0.16}$.
-The lower spectral index can reduce the disagreement with the Northern Tracks sample but requires dedicated analysis.
+We show qualitatively that the lower spectral index from HESE sample can reduce the disagreement with the Northern Tracks sample.
 We also forecast the event spectrum for IceCube-Gen2 for the two different fits.
@@
 The latest 7.5 years (2635 days) high energy starting events (HESE) sample consists of 102 events, of which 60 events have reconstructed deposited energy above 60 TeV Abbasi et al.
-(2021).
+(2021a).
 The IceCube experiment has also observed a Glashow Resonance (GR) event Glashow (1960) in the 4.6 years PeV energy partially contained events (PEPE) sample Aartsen et al.
 (2021a).
-IceCube also maintains a catalog of through-going muon-neutrinos from the northern hemisphere (henceforth called Northern Tracks).
+IceCube also maintains a catalog of muon-like tracks originating from the northern hemisphere (henceforth called Northern Tracks).
 A large fraction of the Northern Tracks are below 100 TeV and originate via atmospheric neutrinos.
 The highest energy events from this sample have astrophysical origin.
-It is abundantly clear that these neutrinos have an extra-galactic origin and are likely to originate in sources capable of accelerating cosmic rays to EeV-scale energies.
+It is abundantly clear that these high-energy neutrinos have an extra-galactic origin and are likely to originate in sources capable of accelerating cosmic rays to EeV-scale energies.
 The neutrinos are produced in subsequent interactions of these ultra-high energy cosmic rays with ambient gas or radiation.
@@
 (2021).
-One such discrepancy is the inferred value of the spectral index in two independent datasets,
-[…]
-We begin with the ansatz that including “dips” in the neutrino flux will result in a lower spectral index from the HESE sample and thus relax the tension.
-Such dips are a consequence of interaction between astrophysical neutrinos and cosmic background.
-If the interaction is through a s-channel process, then the absorption feature is narrowly located around the resonance energy $E_{\nu}^{\rm res}=M_{X}^{2}/2m_{b}$, where $M_{X}$ is the mass of the mediator and $m_{b}$ is the mass of the background particle.
+One such discrepancy is the inferred value of the spectral index in two independent datasets Abbasi et al.
+(2021a, 2022b); Stettner (2020); Abbasi et al.
+(2025),
+$\gamma_{\texttt{\,astro}}^{\text{\,HESE}}=2.87^{+0.20}_{-0.19}\quad\text{and}\quad\gamma_{\texttt{\,astro}}^{\text{\,Nor. Tr.}}=2.37^{+0.09}_{-0.09}\,.$
+The HESE sample selects events (cascades, tracks, and double-cascades) that start inside the instrumented detector volume.
+Any event that deposits its first significant light inside the fiducial volume, without coinciding hits in the outer region, is considered a starting event.
+Such a strong veto on the outer layers ensures that the event is likely to be induced by a neutrino rather than a through-going muon.
+Furthermore, an energy cut-off of 60 TeV is considered to reduce the atmospheric neutrinos from the sample.
+The Northern Tracks sample selects upward-going muon-like tracks that are mostly through-going muons that originate in neutrino interactions outside the instrumented detector volume.
+This sample has low energy-threshold and very good direction reconstruction, which also allows for point-source searches with this dataset.
+The Northern Track events above 100 TeV mostly contain events of astrophysical origin Abbasi et al.
+(2022b, 2025).
+If the two samples originate from the same diffuse flux of astrophysical neutrinos, then the $\sim 2.3\sigma$ discrepancy between the two fits to the spectral index is quite puzzling.
+Interestingly, the Enhanced Starting-Track Event Selection (ESTES) sample, which selects muon-like tracks that start inside the detector irrespective of their direction, reports an intermediate value for the spectral index Abbasi et al.
+(2024),
+$\gamma^{\text{ESTES}}_{\texttt{astro}}=2.58^{+0.10}_{-0.09}.$
+We refer the reader to Fig. 7 in Ref. Abbasi et al. (2025) for a compendium of fits to the parameters of astrophysical neutrino flux from various IceCube datasets.
+The primary aim of this paper is to answer– How are these fits affected if there are dips in the astrphysical neutrino spectrum?
+Such dips are a consequence of interaction between astrophysical neutrinos and a cosmic background.
+If the interaction is through an $s$-channel process, then the absorption feature is narrowly located around the resonance energy $E_{\nu}^{\rm res}=M_{X}^{2}/2m_{b}$, where $M_{X}$ is the mass of the mediator and $m_{b}$ is the mass of the background particle.
 If the interaction is $t$-channel, then a broad spectrum feature is expected Mohanty et al.
@@
 Chauhan and Mohanty (2018) where the dip is caused by interactions between astrophysical neutrinos and a cosmic sterile-neutrino background.
-Figure 1: The test statistic (see Equation 5) for the single power-law fit of 7.5 years HESE sample Abbasi et al.
-(2021) is shown in blue.
-In orange, we show the test statistic after including absorption features and Glashow resonance.
+Figure 1: The test statistic (see Equation 7) for the single power-law fit of 7.5 years HESE sample and Glashow resonance Abbasi et al.
+(2021a) is shown in blue.
+In orange, we show the test statistic after including absorption features.
 The horizontal grey dashed lines are used to denote the 68% C.L.
@@
 intervals.
-We have shown the results from the 11-year Northern Tracks sample Stettner (2020) in green.
-The light sterile neutrino is a well-motivated extension of the standard model and has been the subject of investigation in a number of experiments.
-Currently, there is evidence in support of  Anselmann et al.
-(1995); Hampel et al.
-(1998); Aguilar-Arevalo et al.
-(2009, 2010, 2001); Mention et al.
-(2011); Aguilar-Arevalo et al.
-(2018); Serebrov et al.
-(2021) as well as against Abratenko et al.
-(2023) the sterile neutrino.
-For a recent review, we refer to Ref.
-Dasgupta and Kopp (2021).
-The presence of a fully thermalized sterile neutrino is in conflict with the measurement of $N_{\text{eff}}$ from cosmological probes Hamann et al.
-(2011); Cooke et al.
-(2018); Aghanim et al.
-(2020).
-If the sterile neutrino has self-interactions via a MeV scale vector mediator, their recoupling can be delayed, and the $N_{\text{eff}}$ constraints can be avoided Hannestad et al.
-(2014); Dasgupta and Kopp (2014).
-The case for pseudo-scalar mediator is studied in Ref.
-Archidiacono et al.
-(2015).
-Despite the early success, the self-interacting sterile neutrino scenario has been tested against many cosmological observations Saviano et al.
-(2014); Mirizzi et al.
-(2015); Forastieri et al.
-(2017); Choi et al.
-(2018); Song et al.
-(2018).
-It was shown in Ref.
-Chu et al.
-(2018) that the basic scenario is in conflict with bounds on the sum of neutrino mass, and some resolutions have been provided as well.
-One must note that these results depend on the choice of background cosmology as $\Lambda$CDM, and the bounds will change if other scenarios are considered Roy Choudhury and Naskar (2019); Roy Choudhury and Choubey (2018); Roy Choudhury and Hannestad (2020).
-Recent analyses of large-scale structure data have also indicated the preference for self-interacting neutrinos He et al.
-(2024); Camarena et al.
-(2023).
-With this as motivation, we ask if the dip due to a cosmic sterile-neutrino background can be probed in IceCube and whether the inclusion of these features leads to an appreciable decrease in the inferred spectral index.
-The main result of this paper is shown in Figure 1.
-We find that the systematic inclusion of the dips results in,
+We have shown the results from the 11-year Northern Tracks sample Abbasi et al.
+(2022b) in green.
 […]
+$\gamma_{\texttt{\,astro}}^{\text{\,HESE,\,This Work}}=2.60^{+0.20}_{-0.19}$
 which is in better agreement with the Northern Tracks sample.
-As discussed below, the errors are underestimated due to the limitations of our analysis.
-For a quantitative assessment of the relaxation in the tension between the two datasets, one must evaluate the impact of dips in the spectrum on the Northern Tracks sample.
-However, this exercise is beyond the scope of our paper.
-We have shown that an ‘SPL with dips’ is a viable fit to the present HESE sample and results in a lower spectral index than an SPL-only explanation.
-One must note that our conclusion also extends to other origins of the dip(s).
+As discussed below, our errors are underestimated due to the limitations of our analysis.
+Using a simplified approach, we also show that the inclusion of these dips does not change the fits to the Northern Tracks significantly.
+However, for a quantitative assessment of the relaxation in the tension between the two datasets, a detailed Monte Carlo simulation, along with consistent treatment of backgrounds is required, which is beyond the scope of this work.
+We have shown in detail that an ‘SPL-with-dips’ is a viable fit to the HESE sample and results in a harder spectral index than an SPL-only fit.
+Our conclusions are also applicable to other origins of the dip(s), for example, non-standard interactions of active neutrinos.
 The paper is organized as follows.
-In Section II, we discuss the model, propagation of astrophysical neutrinos, our estimate of the HESE event rate, and the likelihood.
+In Section II, we discuss the methodology followed in the paper.
+We discuss the model and constraints, the propagation of astrophysical neutrinos, our estimate of the event rates the three datasets (HESE, Northern Tracks, and ESTES), and our likelihood analysis.
 In Section III, we provide the main results and end with a short summary in Section IV.
 II Model description and methods
-In this work, we follow the model outlined in Ref.
-Dasgupta and Kopp (2014), which is briefly described below.
+II.1 Lagrangian and constraints
+In this work, we follow the model of self-interacting sterile neutrino outlined in Ref.
+Dasgupta and Kopp (2014).
 The sterile neutrinos are assumed to have self-interactions mediated by a new MeV-scale vector boson.
 The interaction Lagrangian is given by,
-$-\mathcal{L}_{\rm int}=g_{X}\bar{\nu}_{s}\gamma_{\mu}\frac{1}{2}(1-\gamma_{5})\nu_{s}X^{\mu}\,,$ (2)
+$\mathcal{L}_{\rm int}=g_{X}\bar{\nu}_{s}\gamma_{\mu}\frac{1}{2}(1-\gamma_{5})\nu_{s}X^{\mu}\,$, (2)
 where $X^{\mu}$ is the mediator and $g_{X}$ is the self-interaction coupling constant.
@@
 (2015).
+At energy scales lower than the mediator mass ($M_{X}$), the interaction can be described by a Fermi-like contact interaction with an effective coupling given by $G_{X}=\frac{\sqrt{2}}{8}\frac{g_{X}^{2}}{M_{X}^{2}}$.
+Cosmological observations like big bang nucleosynthesis (BBN) and cosmic micorwave backgound (CMB) impose stringent constraints on this effective coupling Song et al.
+(2018).
+It was shown in Ref.
+Song et al.
+(2018) that small values of the effective coupling do not sufficiently delay the production and are thus excluded.
+The minimum value of $G_{X}$ required to satisfy the BBN constraints at $95\%$ C.L.
+is $G_{X}\geq 10^{4}G_{F}$.
+For the mediator mass range relevant to this work, i.e., $M_{X}\sim 10-50$ MeV, $g_{X}>0.01$ is required to satisfy the BBN constraints.
+We have shown resulting parameter space for both fixed-$g_{X}$ and fixed-$G_{X}$.
 The mass-eigenstates and the flavor-eigenstates are related by a $4\times 4$ rotation matrix characterized by six mixing angles.
@@
 In this simplified model, we have five New Physics parameters – three mixing angles ($\theta_{14}\,,\theta_{24}\,,\theta_{34}$), mass of the sterile neutrino ($m_{s}$), and mass of the mediator ($M_{X}$).
-In a UV-complete model, additional degrees of freedom may be required for generating the mixing and mass of the gauge boson and for anomaly cancellation.
-The alternative model with a pseudoscalar mediator is cosmologically more viable Dasgupta and Kopp (2021), but has additional constraints from meson decays Fiorillo et al.
+In a UV-complete model, additional degrees of freedom may be required for generating the mixing and mass of the gauge boson, and for anomaly cancellation.
+Constraints from cosmology– The self-interactions of the sterile neutrino delay their production and hence the constraints from BBN on $N_{\rm eff}$ can be avoided.
+However, the stringest constraints on the model come from the cosmological inference of the sum of neutrino masses Chu et al.
+(2018).
+One of the possible ways to avoid this constraint is to assume additional light sterile neutrinos charged under the new gauge symmetry Chu et al.
+(2018).
+In this scenario, when these sterile neutrinos recouple, the energy is redistributed among all new species which weakens the constraints on the sum of neutrino mass.
+In such a scenario, the number density of the cosmological eV-scale sterile neutrino background will be reduced by a factor proportional to the number of additional species.
+This can be compensated by a small increase in the coupling (as cross-section scales as $g_{X}^{4}$).
+This possibility was considered in Ref.
+Chauhan and Mohanty (2018).
+The other possibility is to consider a fast decay of the sterile neutrino to active neutrinos Chu et al.
+(2018).
+As these sterile neutrinos decay after the decoupling, this would result in a new non-thermal distribution of active neutrinos ($f(p)\sim\delta(p-m_{s}/(2\,\text{or}\,3))$) and this may relax the neutrino mass bound Oldengott et al.
+(2019).
+The scattering of the cosmogeic neutrinos with the new component would result in similar “dips”, however this intersting possibility requires a separate and dedicated analysis.
+Laboratory constraints– In Ref.
+Fiorillo et al.
+(2020a), the constrains on self-interacting sterile neutrino from meson decays was evaluated assuming pseudo-scalar interactions.
+We expect that similar bounds will exist on vector mediators, however, we do not evlauate them.
+These bounds arise from measurements of leptonic decays of charged mesons ($M\rightarrow\ell+\nu_{\ell}$) which get additional contribution from self-interactions of the form $M\rightarrow\ell+\nu_{\ell}+X$ and $M\rightarrow\ell+\nu_{\ell}+\bar{\nu}_{s}+\nu_{s}$.
+The four-body decays are supressed, and dominant constraints arise from three-body decays only Fiorillo et al.
 (2020a).
-We anticipate that similar bounds will exist on vector mediators, but their evaluation is beyond our scope.
-Our main conclusion, i.e., the spectral index is lowered because of the dips, does not depend on the details of the model causing the dips in the spectrum.
+For the light sterile neutrino that is considered in this paper, the bounds on the coupling constant are strong when the new decay modes become kinematically accesible.
+Based on the results of Ref.
+Fiorillo et al.
+(2020a), we expect that our model is strongly constrained for $M_{X}<$ 2 GeV, which is the entire parameter space in consideration.
+These bounds will be sensitive to details of UV completion, e.g., presence of additional light sterile neutrinos and other mediators.
+In this respect, the alternative model with a pseudoscalar mediator allows for selective flavor-dependence of the couplings (unlike gauge interactions that are universal) and the limits can be relaxed Blinov et al.
+(2019).
+Moreover, the pseudoscalar mediator is also cosmologically more viable tooDasgupta and Kopp (2021).
+II.2 Propagation of astrophysical neutrinos
 The introduction of this new interaction results in scattering between astrophysical neutrinos and the cosmic background of active and sterile neutrinos.
@@
 $\frac{\partial\tilde{n}_{i}(t,E_{\nu})}{\partial t}$ $=\frac{\partial}{\partial{E_{\nu}}}\left[H(t)\,E_{\nu}\,\tilde{n}_{i}(t,E_{\nu})\right]+\mathcal{L}_{i}(t,E_{\nu})-\tilde{n}_{i}(t,E_{\nu})\sum_{j}n^{t}_{j}\sigma_{ij}(E_{\nu})$
-[…]
+$+\sum_{j,\,k,\,l}n^{t}_{j}\int_{E_{\nu}}^{\infty}dE_{\nu}^{\prime}\,\tilde{n}_{k}(t,E_{\nu}^{\prime})\frac{d\sigma_{jk\rightarrow il}}{dE_{\nu}}(E_{\nu}^{\prime},E_{\nu})\,.$ (3)
 In this equation, the first term on the right-hand side accounts for the cosmological expansion.
@@
 (2021), the source term is taken to be a single power law with normalization proportional to the star-formation rate.
-The third term accounts for the interaction with background neutrinos, and the last term accounts for the regeneration of lower-energy neutrinos after interactions.
+The third term accounts for the interaction with background neutrinos which depends on the assumed number density of sterile neutrinos today.
+The sterile neutrinos are produced after recoupling from the population of active neutrinos.
+One can do detailed analysis by solving the 4-flavor Boltzmann transport equation (similar to Refs.
+Song et al.
+(2018); Mazumdar et al.
+(2020)) to obtain the exact number density of the sterile neutrinos.
+However, we will assume flavor equilibrium for simplicity and assume that the number density of sterile neutrinos ($n_{s}$) can be approximated as $n_{s}=(1/4)\times n_{a}$, where $n_{a}=2\times 56\,(1+z)^{3}$ is the number density of active neutrinos per flavor in the standard scenario.
+The last term accounts for the regeneration of lower-energy neutrinos after interactions.
 These equations have been solved using the publicly available code nuSIprop Esteban et al.
@@
 Chauhan and Mohanty (2018) where the source term was ad-hoc, and regeneration was ignored.
-To estimate the rate of high-energy starting event (HESE) in IceCube, we use the Monte Carlo sample of simulated events [^1] https://github.com/icecube/HESE-7-year-data-release/ from Ref.
+II.3 High Energy Starting Events (HESE)
+To estimate spectrum of HESE in IceCube, we use the Monte Carlo sample of simulated events [^1] https://github.com/icecube/HESE-7-year-data-release/ from Ref.
 Abbasi et al.
-(2021).
+(2021a).
 The input parameters used to calculate the HESE event spectrum are tabulated in Table 1, and a description of these parameters can be found in Ref.
 Abbasi et al.
-(2021).
+(2021a).
 Of these, we only consider the astrophysical neutrino spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) as variables and fix other parameters to their typical values.
@@
 As we also scan over the mixing angles, the reweighting accounts for the flavor dependence of the interactions.
-Table 1: The set of physical and nuisance parameters that determine the HESE event rate are tabulated.
+The IceCube collaboration uses the effective likelihood defined in Ref.
+Argüelles et al.
+(2019) for estimating the physics parameters and their confidence and credible intervals Abbasi et al.
+(2021a).
+We follow their analysis closely, and present our results in the frequentist statistical methodology.
+For completeness, the likelihood analysis is summarised here and we refer the reader to Ref.
+Abbasi et al.
+(2021a) and Ref.
+Argüelles et al.
+(2019) for additional details.
+To handle the large MC statistical uncertainty due to small sample size in some bins, Ref.
+Argüelles et al.
+(2019) introduces a modified Poisson likelihood function (called effective likelihood),
+$\mathcal{L}_{\text{eff}}(\mu,\sigma,k)=\left(\dfrac{\mu}{\sigma^{2}}\right)^{\frac{\mu^{2}}{\sigma^{2}}+1}\,\Gamma\left(k+\dfrac{\mu^{2}}{\sigma^{2}}+1\right)\left[k!\left(1+\dfrac{\mu}{\sigma^{2}}\right)^{k+\frac{\mu^{2}}{\sigma^{2}}+1}\Gamma\left(\dfrac{\mu^{2}}{\sigma^{2}}+1\right)\right]^{-1},$ (4)
+where $k$ is the observed number of events, $\mu$ is the expected number of events, $\sigma$ is the MC statistical uncertainty for the bin. The likelihood function is given by,
+$\mathcal{L}({\bm{\theta,\eta}})=\prod_{j}^{N_{\text{bins}}}\,\mathcal{L}_{\text{eff}}(\mu_{j}({\bm{\theta,\eta}}),\sigma_{j}({\bm{\theta,\eta}}),k_{j})$ (5)
+where $N_{\text{bins}}$ is the number of bins, $\bm{\theta}$ is the set of parameters of interest, and $\bm{\eta}$ is the set of nuisance parameters.
+Note that, for a given analysis the set of nuisance parameters also includes physical parameters that are not being examined.
+In Ref.
+Abbasi et al.
+(2021a), the data and the MC events are first separated by their morphology (tracks, cascades, double-cascades) and then binned further in two observables.
+The tracks and cascades are binned in their reconstructed energy and reconstructed zenith angle, whereas the double cascades are binned in reconstructed energy and their reconstructed separation between cascades.
+The bin widths are chosen to be comparable to the detector resolution and there are 840 bins in observable quantities.
+These bins contain contributions from atmospheric as well as astrophysical events.
+Table 1: The set of physical and nuisance parameters that determine the HESE spectrum are tabulated.
 We have shown the best-fit parameters for single power law (SPL) obtained in Ref.
 Abbasi et al.
-(2021) and the values/range of parameters used in this work.
+(2021a) and the values/range of parameters used in this work.
 For a description of these variables, see TABLE VI.1 in Ref.
 Abbasi et al.
-(2021)
+(2021a)
 Parameter SPL best-fit This Work
@@
 $\Phi_{\texttt{muon}}$ 1.187 1.00
-$\Delta\gamma$0.053 -0.050
+$\Delta\gamma$ -0.053 -0.050
 $\nu/\overline{\nu}$ 0.998 1.000
@@
 $\varepsilon_{\rm dom}$ 0.952 0.990
-$\varepsilon_{\rm h.o.}$0.055 0.000
+$\varepsilon_{\rm h.o.}$ -0.055 0.000
 $K/\pi$ 1.000 1.000
@@
 Abbasi et al.
-(2021) closely and present our results in the frequentist statistical methodology.
-We report the new physics parameters that are most likely to explain the HESE data and compare it with the single power-law.
-The best-fit parameters and the errors are obtained using the profile likelihood,
-~$\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})=\max_{\bm{\eta}}\mathcal{L}(\bm{\theta,\eta})\cdot\Pi(\bm{\theta,\eta})$ (4)
-where $\bm{\theta}$ is the set of parameters of interest and $\bm{\eta}$ is the set of nuisance parameters.
-The likelihood function used is the effective likelihood defined in Ref.
-Argüelles et al.
-(2019).
-The constraint $\Pi(\bm{\theta,\eta})$ is constant for parameters for which no external data is available.
+(2021a) closely to estimate the new physics parameters that are most likely to explain the HESE data, and compare it with the single power-law.
+The dips in the astrophysical neutrino spectrum will not affect the events from atmospheric neutrinos and muons.
+Moreover, the atmospheric contribution to the starting events in IceCube is primarily determined be the low-energy events and the zenith-angle distributions.
+Hence, in our analysis we will fix the parameters associated with atmospheric events to their nomianal values [^2] We have verified that keeping the atmospheric parameters free does not change the results significantly.
+This is primarily because the atmospheric parameters are determined by the lower energy events and their zenith-angle distributions., and only analyse astrophysical parameters (i.e., $\gamma_{\texttt{astro}}$ and $\Phi_{\texttt{astro}}$).
+Due to this approximation, the uncertainities are underestimated and a comparison in presented in Appendix A.
+The best-fit parameters and the confidence intervals are obtained using the profile likelihood,
+$\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})=\max_{\bm{\eta}}\mathcal{L}(\bm{\theta,\eta})\cdot\Pi(\bm{\theta,\eta})$ (6)
+where $\Pi(\bm{\theta,\eta})$ denote the constraints on parameters.
+For parameters for which no external data is available, $\Pi$ is constant.
 We have used the Python implementation of SLSQP algorithm to minimize the negative of the log of the function and thus obtain the best-fit point $\hat{\bm{\theta}}$.
 The model parameter test statistic is defined as,
-$\text{TS}(\bm{\theta})=-2\,\log\,\left(\dfrac{\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})}{\tilde{\mathcal{L}}^{\texttt{profile}}(\hat{\bm{\theta}})}\right).$ (5)
+[…]
 The results are presented using Wilks’ theorem with appropriate degrees of freedom Wilks (1938).
-To validate our method, we first evaluate the 68 % C.L.
-and 90 % C.L.
-intervals for single-power law and compare them with the contours in Ref.
+II.4 Glashow Resonance
+The HESE sample only includes events for which the interaction vertex is inside the restricted fiducial volume.
+To enhance the sensitivity to Glashow resonance (GR) events, a larger fiducial volume is required, which is achieved by looking at PeV energy partially contained events (PEPE).
+The Monte Carlo simulation provided with the HESE data release can be utilised to predict the number of GR events in the PEPE sample by a simple rescaling.
+Due to the larger effective area, the expected event rate for the PEPE selection is about twice that of the HESE selection Aartsen et al.
+(2021a).
+However, the livetime of PEPE sample is only 4.6 years as opposed to HESE for which the livetime is 7.22 years.
+If $\lambda_{\text{HESE}}$ is the event rate in the GR energy bin (i.e., 5.6 PeV to 7.7 PeV ), obtained from the simulation, then one can estimate the expected number of events in the PEPE sample as,
+$\lambda_{\text{PEPE}}(\bm{\theta,\eta})=\lambda_{\text{HESE}}(\bm{\theta,\eta})\times 2\times(4.6/7.22)$ (8)
+We have verified that the estimates obtained using this method agree very well with the Extended Data Table 1 in Ref.
+Aartsen et al.
+(2021a).
+As only one event has been observed in this bin, we define the likelihood assuming Poisson statistics as,
+$\mathcal{L}_{\text{GR}}(\bm{\theta,\eta})\approx\lambda_{\text{PEPE}}(\bm{\theta,\eta})\,e^{-\lambda_{\text{PEPE}}(\bm{\theta,\eta})}.$ (9)
+The likelihood function, i.e., $\mathcal{L}(\bm{\theta,\eta})$ in Equation 5, is multiplied by $\mathcal{L}_{\text{GR}}(\bm{\theta,\eta})$ to obtain the combined fit.
+The impact of including the GR event in our analysis is shown in Appendix A.
+II.5 Northern Tracks
+The Northern Tracks analysis by IceCube focusses on measuring the flux of high-energy $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ from atmospheric and astrophysical sources.
+The event selection criterion samples muon-like [^3] We refer to both $\mu^{-}$ and $\mu^{+}$ as “muons” tracks originating from the Northern celestial hemisphere, which significantly reduces the background of atmospheric muons and results in high-purity neutrino-induced events.
+The sample contains mostly through-going muon tracks and a small contribution from starting tracks.
+Unlike the HESE sample, a detailed MC simulation and likelihood analysis for Northern Tracks is not publicly available.
+In order to estimate the impact of dips in the neutrino spectrum on Northern Tracks, we use a “theorist’s approach” Kistler and Beacom (2006); Laha et al.
+(2013); Gaisser et al.
+(2016) to calculate the event rates.
+The measurable muon spectrum for Northern Tracks is given by,
+$\dfrac{dN_{\text{NT}}}{dE_{\mu}^{\text{true}}}=\dfrac{dN_{\text{thru. tr.}}}{dE_{\mu}^{\text{true}}}+\dfrac{dN_{\text{st. tr.}}}{dE_{\mu}^{\text{true}}}$ (10)
+where the first term is the contribution from the through-going tracks that originate outside the detector volume, and the second term denotes the contribution of the starting tracks for which the primary interaction vertex is inside the detector volume.
+We have used $E_{\mu}^{\text{true}}$ to denote the true energy of the muon which is different from the reconstructed muon energy.
+The muon spectrum from the through-going events is calculated using,
+$\dfrac{dN_{\text{thru.
+tr.}}}{dE_{\mu}^{\text{true}}}=\sum_{\nu_{\mu},\bar{\nu}_{\mu}}\dfrac{T\,N_{A}}{\alpha+\beta E_{\mu}^{\text{true}}}\int d\Omega\int_{E_{\mu}^{\text{true}}}^{E_{\nu}^{\text{max}}(\theta_{z})}dE_{\nu}\,A(E_{\nu},\theta)e^{-\uptau(E_{\nu},\theta_{z})}\,\dfrac{d\Phi_{\nu}}{dE_{\nu}}\int_{0}^{1-E_{\mu}^{\text{true}}/E_{\nu}}dy\,\dfrac{d\sigma_{\nu}(E_{\nu},y)}{dy},$ (11)
+[…]
+The evaluation of Equation 11 is not only computationally expensive but also require input from IceCube about effective areas.
+We can simplify the expression by first assuming that $A(E_{\nu},\theta)\approx A=1\,\text{km}^{2}$.
+We also assume that $E_{\nu}^{\text{max}}(\theta_{z})\rightarrow\infty$.
+As a result, the expression can be simplified as,
+$\dfrac{dN_{\text{thru.
+tr.}}}{dE_{\mu}^{\text{true}}}=\dfrac{2\pi\,T\,N_{A}\,A}{\alpha+\beta E_{\mu}^{\text{true}}}\int_{E_{\mu}^{\text{true}}}^{\infty}dE_{\nu}\,\mathcal{S}_{\text{NT}}(E_{\nu})\,\tilde{\sigma}(E_{\mu}^{\text{true}},E_{\nu})\,\dfrac{d\Phi_{\nu+\bar{\nu}}}{dE_{\nu}}$ (12)
+where,
+$\tilde{\sigma}(E_{\mu}^{\text{true}},E_{\nu})=\int_{0}^{1-E_{\mu}^{\text{true}}/E_{\nu}}dy\,\dfrac{1}{2}\,\left[\dfrac{d\sigma_{\nu}(E_{\nu},y)}{dy}+\dfrac{d\sigma_{\bar{\nu}}(E_{\nu},y)}{dy}\right]$ (13)
+and the average attenuation (or shadow) factor for Northern Tracks is,
+$\mathcal{S}_{\text{NT}}(E_{\nu})=\int_{-1}^{0}d(\cos\theta_{z})\,\dfrac{1}{2}\,\left[e^{-\uptau(E_{\nu},\theta_{z})}+e^{-\uptau(E_{\bar{\nu}},\theta_{z})}\right].$ (14)
+Note that in Equation 12, the flux accounts for both $\nu_{\mu}$ and $\bar{\nu}_{\mu}$.
+We have neglected the contributions from leptonic interaction channels such as $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{\mu}+\mu^{-}$ as their cross-sections are significanly smaller than DIS.
+For the starting-tracks contribution, we have followed Ref. Mena et al. (2014) and use the same approximations to obtain,
+$\dfrac{dN_{\text{st.
+tr.}}}{dE_{\mu}^{\text{true}}}=2\pi\,T\,N_{A}\,\rho\,V\int_{0}^{\infty}dE_{\nu}\,\mathcal{S}_{\text{NT}}(E_{\nu})\,\dfrac{d\sigma(E_{\mu},E_{\nu})}{dE_{\mu}}\,\dfrac{d\Phi_{\nu+\bar{\nu}}}{dE_{\nu}}$ (15)
+where the differential cross-section $d\sigma/dE_{\mu}$ is averaged over neutrino and anti-neutrino, $\rho=0.9167$ g/cm3 is the density of ice, and we take $V=1\,\text{km}^{3}$.
+Lastly, we need to account for the detector resolution.
+The true energy deposited by the travelling muon ($E_{\mu}^{\text{true}}$) and the proxy muon energy that is reconstructed do not usually have a one-to-one mapping.
+We model this uncertainty as a Gaussian, and the measured muon spectrum is given by Mena et al.
+(2014),
+$\dfrac{dN}{dE_{\mu}}=\int_{0}^{\infty}dE_{\mu}^{\text{true}}\,\dfrac{dN}{dE_{\mu}^{\text{true}}}\,\mathcal{G}\left(E_{\mu},E_{\mu}^{\text{true}},\sigma(E_{\mu}^{\text{true}})\right)$ (16)
+where we have considered $\sigma(E_{\mu}^{\text{true}})=0.125\,E_{\mu}^{\text{true}}$ in our analysis, similar to Ref.
+Mena et al.
+(2014).
+We understand that the starting track events have a better energy resolution than the through-going tracks, however, we do not make this difference.
+We use Equation 12 and Equation 16 to estimate the contribution of astrophysical neutrinos to the Northern Tracks.
+For atmospheric contribution, we use the results in Ref.
 Abbasi et al.
 (2025).
-We only scan over the astrophysical flux parameters and keep other parameters fixed to their best-fit values.
-We find that on expected lines, the best-fit point agrees very well, and the errors are underestimated.
-The HESE sample only includes events for which the interaction vertex is inside the restricted fiducial volume.
-To enhance the sensitivity to Glashow resonance (GR) events, a larger fiducial volume is required, which is achieved by looking at PeV energy partially contained events (PEPE).
-As the Monte Carlo event sample contains the GR, we use a simple rescaling for exposure to estimate the GR event rate.
-We find that our estimates agree with the Extended Data Table 1 in Ref.
-Aartsen et al.
-(2021a).
-The likelihood function, i.e., $\mathcal{L}(\bm{\theta,\eta})$ in Equation 4, is modified to include GR event for which we assume a simple Poisson likelihood.
+The observed data for Northern tracks is provided in Ref.Abbasi et al.
+(2022b) , which we use for our likelihood analysis.
+The IceCube collaboration uses both energy distribution and zenith-angle distribution of the observed events to calculate the likelihood.
+The atmospheric neutrino contribution to the Northern Tracks dominates at lower energies and has a very strong and characteristic zenith-angle dependence.
+In this work, we only consider the muon-energy bins above 100 TeV to minimize the effect of atmospheric neutrinos.
+We do not consider binning in zenith-angle as the atrophysical neutrinos have almost flat distribution.
+We define the likelihood assuming Poisson distribution as,
+$\mathcal{L}_{i}(\bm{\theta},\bm{\eta})=\frac{\lambda_{i}(\bm{\theta},\bm{\eta})^{k}e^{-\lambda_{i}(\bm{\theta},\bm{\eta})}}{k!}$ (17)
+where $\lambda_{i}(\bm{\theta},\bm{\eta})$ is the expected number of events in the $i$-th bin obtained by integrating Equation 16, and $k$ is the number of number of observed events in the bin.
+In our analysis, we only consider the bins for which events are recorded and we do not consider the zenith-angle distribution of events.
+The likelihood, profile likelihood, and the test statistic are defined similar to Equation 5, Equation 6, and Equation 7, respectively.
+II.6 Enhanced Starting-Track Event Selection (ESTES)
+The Enhanced Starting-Track Event Selection (ESTES) in IceCube is a event selection criterion that focuses on identifying $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ interactions that originate inside the detector volume from all directions (northern sky and southern sky) Silva (2021); Silva and Mancina (2020); Mancina and Silva (2020, 2021); Abbasi et al.
+(2021b); Silva et al.
+(2023); Silva (2023); Abbasi et al.
+(2024).
+By employing a dynamical veto volume and neutrino self-veto, the backgrounds from atmospheric muons and atmospheric neutrinos can be efficiently rejected, especially from the southern sky.
+This allows IceCube to probe for astrophysical muon neutrinos down to 1 TeV.
+ESTES events benefit from good directional resolution and improved neutrino energy reconstruction compared to other event selections, and have enchanced sensitivity to southern-sky neutrino sources, including the Galactic plane.
+These event can also be used to study the diffuse flux of neutrinos.
+Assuming a single power-law, the best fit spectral index for astrophysical neutrinos is,
+$\gamma^{\text{ESTES}}_{\texttt{astro}}=2.58^{+0.10}_{-0.09},$ (18)
+which lies between the fits obtained from HESE and Northern Tracks.
+In order to estimate the impact of dips on the ESTES data, we again adopt a theorist’s approach to event rates.
+The muon spectrum for starting-tracks can be obtained using Equation 15 and Equation 16.
+As the ESTES sample looks at events from all directions, the shadow factor can be approximated as,
+$\mathcal{S}_{\text{ESTES}}(E_{\nu})=1+\mathcal{S}_{\text{NT}}(E_{\nu}),$ (19)
+assuming that the events from the southern-sky are not attenuated. Our approximation for the energy resolution should work for ESTES data as wellAbbasi et al. (2024).
+Similar to Northern Tracks, the IceCube collaboration uses both muon-energy and zenith-angle distribution for their statistical analysis.
+In this work, we will only only consider the muon-energy bins above 70 TeV for the likelihood analysis of the ESTES data.
+The bin likelihood, total likelihood, profile likelihood, and the test statistic are defined similar to Equation 17, Equation 5, Equation 6, and Equation 7, respectively.
 III Searching for dips in spectrum
+III.1 HESE and GR Analysis
 In this section, we look at the range of new physics parameters that lead to significant dips in the HESE sample and evaluate their impact on the spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) of the diffuse astrophysical flux.
@@
 region is mostly unaffected.
-Table 2: The best-fit values of the parameters of interest for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the result from our analysis is tabulated.
+Figure 2: Left: Scan of the profile likelihood (see Equation 7) for the two new physics parameters: mass of the sterile neutrino ($m_{s}$) and mass of the vector-boson mediator ($M_{X}$), is shown in orange.
+The darker (lighter) region denotes the 68% C.L.
+(90 % C.L.) interval.
+Note that for each point, the mixing angles are fixed to their best-fit values, and all other parameters are optimized.
+In this plot, we have fixed $g_{X}=0.1$.
+We also show two approximate curves that result in dips at 300 TeV and 800 TeV due to interaction with a cosmic sterile neutrino background.
+The region labeled as NSI (pink) indicated the parameter space for which dips occur due to interaction with the active neutrino background only and the dip due to sterile neutrino is outside the HESE analysis bound.
+Right: Same as the left figure but for $G_{X}=10^{6}\,G_{F}$
+Table 2: The best-fit values of the parameters of interest for single power law (SPL) obtained in Ref. Abbasi et al. (2021a) and the result from our HESE+GR analysis is tabulated.
 Parameter SPL best-fit This Work
@@
 $\theta_{34}$ 0.41
-Figure 2: Scan of the profile likelihood (see Equation 5) for the two new physics parameters: mass of the sterile neutrino ($m_{s}$) and mass of the vector boson mediator ($M_{X}$), is shown in orange.
-The darker (lighter) region denotes the 68% C.L.
-(90 % C.L.) interval.
-Note that for each point, the mixing angles are fixed to their best-fit values, and all other parameters are optimized.
-We also show two approximate curves that result in dips at 300 TeV and 800 TeV due to interaction with a cosmic sterile neutrino background.
-The region labeled as NSI (pink) indicated the parameter space for which dips occur due to interaction with the active neutrino background only and the dip due to sterile neutrino is outside the HESE analysis bound.
-We first look at the space of the new physics parameters – sterile neutrino mass ($m_{s}$) and the vector mediator mass ($M_{X}$), and present our results in Figure 2.
-As expected [^2] The Figure 2 should be compared with Fig.
-2 in Chauhan and Mohanty (2018), we find a large degenerate parameter space that results in dips in the event spectrum.
-We also find an island for $M_{X}=5-10$ MeV, which is largely independent of the sterile neutrino mass.
-In this region, the dip due to sterile neutrino is below the 60 TeV threshold of the HESE sample, and there is a dip between 300-800 TeV due to the heavier active neutrino.
-This is similar to the case of neutrino secret interaction ($\nu$SI), which is considered in Ref.
-Ng and Beacom (2014).
-We also highlight the $m_{s}=2.6$ eV that is hinted in recent results Serebrov et al.
-(2021); Abbasi et al.
-(2022b).
-Figure 3: Scan of the profile likelihood for the two new physics parameters: astrophysical normalization and spectral index.
-The darker (lighter) regions show the 68 % C.L.
-(90 % C.L.) intervals.
-The region in blue is obtained using single power law, and the region in orange is obtained by including Glashow Resonance and dips due to sterile neutrino background.
-Note that other physical parameters are optimized for each scan point.
-As we do not optimize over all of the parameters (see text for details), the intervals are slightly smaller than the ones reported in Ref.
-Abbasi et al.
-(2021).
-Figure 4: Left: We show the best-fit SPL flux from Ref.
-Abbasi et al.
-(2021) (blue) and the best-fit flux from our analysis (orange).
-The binned reconstructed flux from the HESE sample Abbasi et al.
-(2021) and PEPE sample is shown in black and red, respectively.
-Right: The deposited energy spectrum of HESE events (black) is shown with our event prediction using the best-fit parameters for SPL with dips (orange) and SPL-only (blue).
-The background from atmospheric neutrinos and muons is the same for both curves.
-The shaded grey region represents the 60 TeV threshold below which the backgrounds dominate the event rate.
+[…]
 We now look at the space of astrophysical neutrino flux parameters – spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$).
@@
 This is one of the key results of the paper.
-At present, with only 7.5 years of data, there is a large overlap between the preferred regions for SPL and SPL with dips.
+At present, with only 7.5 years of data, there is a large overlap between the preferred regions for SPL and SPL-with-dips.
 However, the trend is clearly noticeable.
@@
 We also look at single parameter profile likelihood, and the results have been presented in Figure 1.
-It is clear that the inclusion of absorption features leads to a lowering of the spectral index and thus reduces the tension in the HESE and Northern Tracks samples.
-A future systematic study would also include the impact of absorption on the Northern Tracks sample.
-In order to highlight the difference between the SPL best-fit and the best-fit parameters for SPL with dips, we show the neutrino flux prediction overlaid with the reconstructed flux from the IceCube collaboration in Figure 4.
+It is clear that due to the inclusion of dips in the spectrum, the fits to HESE and GR data prefer a harder spectral index.
+In order to highlight the difference between the SPL best-fit and the best-fit parameters for SPL-with-dips, we show the neutrino flux prediction overlaid with the reconstructed flux from the IceCube collaboration in Figure 4.
 This plot only indicates of the position of dips in the neutrino spectrum.
@@
 We have shown the predicted event spectrum from the two sets of best-fit parameters in Figure 4.
-As mentioned earlier, we assume that the atmospheric neutrinos are unaffected by the new physics, and the background predictions for both fits are the same.
-In a more detailed analysis, one may consider varying the atmospheric neutrino parameters.
+As mentioned earlier, we have verified that the atmospheric neutrino paramters are unaffected by the new physics, and the background predictions for both fits are statistically indistinguishable.
+In a more detailed analysis, one may consider varying the models for atmospheric neutrinos.
 As there is a large overlap in the 68% C.L.
 interval of the two fits, we choose not to show the uncertainty bands in these figures.
-Figure 5: Event rate prediction for 25 years of exposure for IceCube-Gen2.
-The orange curve represents events assuming the best-fit parameters of this work (see Table 2).
-The blue curve represents events for the SPL best-fit parameters.
-The blue-shaded region shows 68 % C.L.
-errors on the SPL event rate assuming Poisson distribution.
-The grey-shaded region represents the 200 TeV threshold for IceCube-Gen2 Aartsen et al.
-(2021b).
-III.1 Forecast for IceCube-Gen2
+Figure 3: Scan of the profile likelihood for the two astrophysical neutrino flux parameters: normalization scale and spectral index obtained from analysis of HESE and GR data.
+The darker (lighter) regions show the 68 % C.L.
+(90 % C.L.) intervals.
+The region in blue is obtained using single power law, and the region in orange is obtained by including the dips due to sterile neutrino background.
+Note that other physical parameters are optimized for each point.
+As we do not optimize over all of the parameters (see text for details), the intervals are slightly smaller than the ones reported in Ref.
+Abbasi et al.
+(2021a).
+See Appendix A for comparison.
+Figure 4: Left: We show the best-fit SPL flux from Ref.
+Abbasi et al.
+(2021a) (blue) and the best-fit flux from our analysis (orange).
+The binned reconstructed flux from the HESE sample Abbasi et al.
+(2021a) and PEPE Aartsen et al.
+(2021a) sample is shown in black and red, respectively.
+Right: The deposited energy spectrum of HESE events (black) is shown with our event prediction using the best-fit parameters for SPL with dips (orange) and SPL-only (blue).
+The background from atmospheric neutrinos and muons is the same for both curves.
+The shaded grey region represents the 60 TeV threshold below which the backgrounds dominate the event rate.
+III.2 Northern Tracks Analysis
+In Section II.5, we outlined our methodology for estimating the distribution of Northern Track events in IceCube.
+We now compare our predictions with observed IceCube data to extract the astrophysical flux normalization and spectral index.
+We find that our methods systematically underestimate the event rates.
+This is a significant limitation of our analysis, and our best-fit normalization differs substantially from the official IceCube results.
+A further limitation of our analysis is the treatment of the atmospheric backgrounds.
+The fit of atmospheric neutrino parameters are determined by the low-energy data as well as their zenith-angle distribution, and is largely independent of the dips.
+Subsequently, we adopt the atmospheric component as reported by IceCube, and only use the high-energy bins (>100>100\,TeV) for our likelihood analysis.
+Moreover, we do not utilise the zenith-angle distribution in our analysis.
+As a result of this simplified treatment of statistics, the uncertainities on our analysis are over-estimated.
+Therefore, our results should be interpreted in terms of trends and relative comparisons rather than as precise determinations of the flux parameters.
+With these limitations in mind, we examine whether the Northern Track data can discriminate between two representative flux hypotheses, i.e., SPL and SPL-with-dips.
+For the SPL-with-dips analysis, we fix the New Physics parameters to the best-fit values from the HESE-only analysis (see Table 2).
+The results are presented in Figure 5, where we show the test-statistic for the two scenarios as well as our predicted event distributions.
+As can be seen, Northern Tracks are largely insensitive to the dips in the astrophysical neutrino spectrum.
+This insensitivity arises because the observed muons are predominantly through-going, and thus the observed muon energy does not correspond uniquely to the incident neutrino energy [^6] A neutrino with larger energy interacting further from the detector and a neutrino will lesser energy interacting closer to the detector lead to a through-going muon of similar energy..
+As a result, any sharp spectral features are smeared out in the observed muon spectrum, and leave only a modest imprint on the event rates.
+The detector resolution (see Equation 16) leads to additional smearing for both hypotheses equally and is not the dominant cause of this insensitivity.
+As a result, the prediction for spectral index is indistinguishable for the two scenarios (see Figure 5, left).
+This is also evident in our comparison of the event distributions for the the best-fit values for the two hypothesis (see Figure 5, right).
+The event rates are statistically indistinguishable.
+The predicted normalizations for the two fits are slightly different.
+As mentioned earlier, our simplified treatment does not yield the correct normalisation and hence we do not report them.
+Due to the insensitivity of Northern Tracks data to the narrow spectral features, changing the location of the dips (i.e.
+a scan over $M_{X}$ and $m_{s}$) does not yield meaningful insights.
+On the other hand, broad spectral features (e.g., those considered in Fiorillo and Bustamante (2023) and Mohanty et al.
+(2019)) can lead to more pronounced differences and therefore are more testable with Northern Tracks.
+Figure 5: Left: The test statistic (see Equation 7) for the single power-law fit of 9.5 years Northern Tracks sample Abbasi et al.
+(2022b) is shown in blue.
+In green, we show the test statistic after including absorption features using the best-fit New Physics parameters from Table 2.
+The horizontal grey dashed lines are used to denote the 68% C.L.
+and 90% C.L.
+intervals.
+Right: The event rates obtained using Equation 16 for the SPL and SPL+dips are shown in blue and green respectively.
+Note that the normalization for both fits is slightly different.
+The observed data points are taken from Abbasi et al.
+(2022b) and the errors are statistical only.
+The red shaded region is the atmospheric contribution.
+In the grey shaded region, the atmospheric contribution is large and hence excluded from our analysis (see text for details).
+III.3 ESTES Analysis
+In Section II.6, we outlined our methodology for estimating the distribution of ESTES events in IceCube.
+Similar to Northern Tracks analysis, we now compare our predictions with observed IceCube data to extract the astrophysical flux parameters as well as determine the sensitivity to dips in the spectrum.
+We again find that our simple methods systematically underestimate the ESTES event rates when compared to IceCube’s detailed simulations.
+Similar to northern tracks analysis, we fix the atmospheric neutrino background and use only the high energy bins (>70>70\,TeV) in the likelihood.
+Again, the results in this section should be interpreted as trends and relative comparisons rather than as precise determinations.
+Our ESTES analysis is similar to Northern Tracks and the results are presented in Figure 6.
+Since these are track-like events that start inside the detector, there is a better relationship between the measured muon energy and the incident neutrino energy.
+As a result, any narrow spectral features lead to difference in the predicted spectral index (see Figure 6, left).
+The best fit values from our analysis are,
+[…]
+The lowering of the spectral index is comparable to the results from HESE-only analysis.
+We also compare the event distributions for the the best-fit values for the two hypothesis in Figure 5 (right).
+The predicted normalizations for the two fits are different.
+In addition, we also show the event rates obtained by considering dips and the SPL best-fit values for spectral index and normalization for comparison.
+Figure 6: Left: The test statistic (see Equation 7) for the single power-law fit of 10.5 years ESTES sample Abbasi et al.
+(2024) is shown in blue.
+In purple, we show the test statistic after including absorption features using the best-fit New Physics parameters from Table 2.
+The horizontal grey dashed lines are used to denote the 68% C.L.
+and 90% C.L.
+intervals.
+Right: The event rates obtained using Equation 16 for the SPL and SPL+dips are shown in blue and purple respectively.
+Note that the normalization for both fits is different.
+The purple dashed curve is obtained by considering dips with flux parameters from best-fit values of SPL, and shown only for comparison.
+The observed data points are taken from Abbasi et al.
+(2024) and the errors are statistical only.
+The red shaded region is the atmospheric contribution.
+In the grey shaded region, the atmospheric contribution is large and hence excluded from our analysis (see text for details).
+III.4 Forecast for IceCube-Gen2
 IceCube-Gen2 is the planned upgrade of the IceCube neutrino observatory, which is expected to have $\sim$10 times larger exposure to TeV–PeV neutrinos Aartsen et al.
@@
 In this work, we will assume that the effective area is the same for cascades and tracks.
-We have evaluated the event rates using 25 years of IceCube-Gen2 exposure for the best-fit parameters from the SPL analysis and for the best-fit point from this work.
+From the preceeding discussion, it is clear that starting events (HESE or ESTES) are good probes of dips in the spectrum.
+We have evaluated the HESE event rates using 25 years of IceCube-Gen2 exposure for the best-fit parameters from the SPL analysis and for the best-fit New Physics parameters from this work.
 The results are presented in Figure 7.
@@
 A combined analysis will improve the discovery prospects Fiorillo and Bustamante (2023).
+Figure 7: Event rate prediction for 25 years of exposure for IceCube-Gen2.
+The orange curve represents events assuming the best-fit parameters of this work (see Table 2).
+The blue curve represents events for the SPL best-fit parameters.
+The blue-shaded region shows 68 % C.L.
+errors on the SPL event rate assuming Poisson distribution.
+The grey-shaded region represents the 200 TeV threshold for IceCube-Gen2 Aartsen et al.
+(2021b).
 IV Summary and outlook
@@
 It also provides an opportunity to study rare interactions over large distances.
-In this paper, we have taken a close look at the interaction of the astrophysical neutrinos with a hypothetical sterile neutrino background.
+In this paper, we have taken a close look at the interaction of the astrophysical neutrinos with a hypothetical sterile neutrino background from the perspective of various IceCube datasets.
 The light sterile neutrino is well-motivated from short baseline experiments.
@@
 The astrophysical neutrinos can interact with the sterile neutrino background, which leads to dips in the TeV–PeV neutrino energy spectrum.
-If the dips are significant, there is an observable impact on the HESE sample in IceCube.
-We have improved on the previous analysis on multiple fronts.
+If the dips are significant, there is an observable impact on the starting events such as HESE and ESTES.
+The Northern Tracks are not significantly affected as there is a lack of one-to-one correspondence between observed event energy and the incident neutrino energy.
+We have improved on the previous HESE analysis on multiple fronts as well.
 We have taken a realistic source distribution and included the effect of regeneration during the propagation.
@@
 Note that, due to the approximations in our analysis, our errors are underestimated.
-The lower spectral index is in better agreement with the Northern Tracks sample ($\gamma_{\texttt{astro}}=2.28^{+0.08}_{-0.09}$).
-However, a detailed study that accounts for the effect of these absorption features on the Northern Tracks sample is essential and left as future work.
+The lower spectral index is in better agreement with the Northern Tracks sample ($\gamma_{\texttt{astro}}=2.37^{+0.09}_{-0.09}$), which are largely insensitive to the presence of such narrow features in the neutrino spectrum.
+The ESTES sample can prove complementary probe of these dips and can provide information about the flavor structure, if any.
 Currently, with only 7.5 years of HESE data, we find that there is a large overlap of parameter space, and one cannot statistically infer the existence of such dips.
@@
 The authors would also like to thank Subhendra Mohanty and Ivan Esteban for their comments and suggestions on the manuscript.
-PP thanks Ranjan Laha for useful discussion.
+The authors thank Ranjan Laha for useful discussions.
+The authors acknowledge the anonymous referee who motivated a large body of work in this manuscript.
 BC is supported by an OPERA grant and seed grant NFSG/PIL/2023/P3821 from BITS Pilani.
 PP acknowledges the support from the National Post-Doctoral Fellowship by the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (SERB-PDF/2023/002356) and the IOE-IISc fellowship program.
+The authors acknowledge the use of computational resources provided by DIST-FIST Project No.
+SR/FST/PS-1/2017/30.
+Appendix A
+In order to estimate the effects of our approximations on the fit, we evaluate the 68% C.L.
+and 90% C.L.
+intervals for single-power law and compare them with the contours in Ref.
+Abbasi et al.
+(2021a).
+The results are shown in Figure 8.
+We only scan over the astrophysical flux parameters and keep other parameters fixed close to their best-fit values.
+We find that on expected lines, the best-fit point agrees very well, and the errors are underestimated.
+It is also seen that the impact of these approximations lead to a larger change in the normalization and the impact on the spectral index is small.
+Figure 8: The SPL fit to HESE data Abbasi et al.
+(2021a) considering all nuisance parameters (see Table 1) is shown in grey.
+The inner (dark grey) region shows the 68 % C.L.
+interval and the outer (light grey) region shows the 90 % C.L.
+interval.
+The best-fit point is shown as a grey star.
+Our fit to HESE data in which the nuisance parameters are fixed to their nominal values is shown as red curves.
+Our best-fit point for HESE-only analysis is very close to the IceCube best-fit point and not shown here.
+The fit including HESE and Glashow Resonance (GR) is shown in blue.
+The blue star shows the best-fit point of our HESE+GR analysis.
+The impact of including the Glashow Resonance event is also on expected lines.
+In order to get a reasonable expectation for the energy bin, one must have a harder spectrum, and larger values of spectral index are not preferable.
+As one can see in the Figure 8, the fits for smaller values of $\gamma_{\texttt{astro}}$ for both HESE and HESE+GR are similar, and differ for larger values.